Write equivalent equations in the form of inverse functions for
a. x= y+ cos0(that 0 is suppose to have a line through it)
b. cosy= x(to the second power)

Answer by stanbon(75887)

Write equivalent equations in the form of inverse functions for
===============================
a. x= y+ cos0(that 0 is suppose to have a line through it)
cos(theta) = x-y
theta = cos^-1(x-y)
=======================

b. cosy= x(to the second power)
cos(y) = x^2
y = Cos^-1(x^2)
=================
Cheers,
Stan H.